OpenStudy (anonymous):

-3(2x-1)4(5y-2x) Simplify using algebraic expression b combining like terms. Please show step by step.

OpenStudy (anonymous):

Hint: you can't combine like terms until you expand your expression out! In other words, multiply everything together first.

OpenStudy (anonymous):

yes I'm fine with that I just keep getting confused about negatives and positives. Example in this problem is it -4*5y or is it 4*5y ( I forgot to put the - sign in front of the 4 as in the original problem

OpenStudy (anonymous):

Well, if you were distributing just 4(5y-2x), you're multiplying (5y-2x) by a *positive* 4, so you don't have to worry about sign changes: 4(5y-2x) = 20y-8x On the other hand, with -3(2x-1), you're multiplying by a -3 so -3(2x-1) = -3(2x) - (-3)(1) = -6x + 3 Does that make sense?

OpenStudy (anonymous):

-3(2x-1)-4(5y-2x) is how my problem should have read. I'm just confused with the symbol before the 4. is this minus or negative 4. Just when i think i am fine then i confuse myself again.

OpenStudy (anonymous):

Oh! Okay. Thank you for the clarification. So, with the -4, we're going to be multiplying through by a *negative* -4(5y-2x) = -20y + 8x Now, we can add it to your original expression -3(2x-1) -20y + 8x Does that make sense? Another way to think about it would be: -3(2x-1)-4(5y-2x) = -3(2x-1) - [4(5y-2x)] <-- added [ ] So focusing on the 4 first: -3(2x-1) - [20y-8x] And now distribute that negative: -3(2x-1) - 20y + 8x

OpenStudy (anonymous):

Thank you! The bars around the 4 also helped. So the end result is2x+3 -20y?

OpenStudy (anonymous):

You're correct :) and you're welcome!

OpenStudy (anonymous):

=)